### v1

@@ 6 Conclusion
In this work, we have shown the equivalence of two different pinning control protocols, additive pinning and parametric pinning, for oscillatory and periodic weakly coupled systems. While additive pinning is often implemented in engineering applications, parametric pinning seems more difficult to implement in practice. In fact, in experiments, it is easier to add an external input, while the task of acting on the parameters of the systems to be controlled seems more challenging. Nonetheless, in certain systems, such as social systems, it is easier to directly act on the model parameters rather than applying an external input. For instance, in case of an epidemic spreading, one can control a given location, i.e., a node of the network, by implementing certain containment measure. In practice, this would mean acting on the parameters of that given node, i.e., parametric pinning. It is true that we have shown the equivalence of the two protocols only for oscillatory systems and not in general. However, oscillatory behaviors occur also in such social systems and epidemics: for instance, when considering a periodic infectivity parameter, it was shown that the system exhibits a periodic attractor [4]. Therefore, our result is twofold: from a purely theoretical point of view, it deepens our knowledge of the phase reduction method for systems to which a control is applied; on the other hand, it also provides guidance for applications, by formally proving the equivalence of the two control protocols, once certain conditions are satisfied.
@@ Acknowledgements
The authors are grateful to Hiroya Nakao, Mattia Frasca, and Lucia Valentina Gambuzza for discussions. The work of R.M. is supported by a JSPS postdoctoral fellowship, grant 24KF0211. Y.K. acknowledges JSPS KAKENHI JP22K14274 and JST PRESTO JPMJPR24K3 for financial support.

### v3

@@ 6 Conclusion
In this work, we have shown the equivalence of two different pinning control protocols, additive pinning and parametric pinning, for oscillatory and periodic weakly coupled systems. While additive pinning is often implemented in engineering applications, parametric pinning seems more difficult to implement in practice. In fact, in experiments, it is easier to add an external input, while the task of acting on the parameters of the systems to be controlled seems more challenging. Nonetheless, in certain systems, such as social systems, it is easier to directly act on the model parameters rather than applying an external input. For instance, in case of an epidemic spreading, one can control a given location, i.e., a node of the network, by implementing certain containment measures. In practice, this would mean acting on the parameters of that given node, i.e., parametric pinning. It is true that we have shown the equivalence of the two protocols only for oscillatory systems and not in general. However, oscillatory behaviors occur also in such social systems and epidemics: for instance, when considering a periodic infectivity parameter, it was shown that the system exhibits a periodic attractor [4]. Therefore, our result is twofold: from a purely theoretical point of view, it deepens our knowledge of the phase reduction method for systems to which a control is applied; on the other hand, it also provides guidance for applications, by formally proving the equivalence of the two control protocols, once certain conditions are satisfied.
@@ Acknowledgements
The authors are grateful to Hiroya Nakao, Mattia Frasca, and Lucia Valentina Gambuzza for discussions. R.M. acknowledges JSPS KAKENHI 24KF0211 for financial support. Y.K. acknowledges JSPS KAKENHI JP22K14274, and JST PRESTO JPMJPR24K3 for financial support.